Find the center and radius of the circle that passes through M(3,3) and is tangent to the line Y-2x+3 at point N(5,5)


1
Expert's answer
2021-09-09T00:41:25-0400

The line "y=2x+3" does not pass through the point "N(5,5)"


"2(5)+3=13\not=5"

Assume that the problem is:

Find the center and radius of the circle that passes through "M(3,3)" and is tangent to the line "y=2x+3" at point "N(1,5)"


The general equation for a circle is "( x - h )^2 + ( y - k )^2 = r^2," where "C( h, k)" is the center and "r" is the radius.

The circle that passes through "M(3,3)"


"( 3 - h )^2 + (3 - k )^2 = r^2"

The circle that passes through "N(1,5)"


"( 1 - h )^2 + (5 - k )^2 = r^2"


The line segment "CN" is perpendicular to the line "y=2x+3"


"slope_1(slope_2)=-1"

"\dfrac{y_N-y_C}{x_N-x_C}(2)=-1"

"\dfrac{5-k}{1-h}=-\dfrac{1}{2}"

"10-2k=-1+h"

"h=11-2k"

"( 3 - h )^2 + (3 - k )^2 =( 1 - h )^2 + (5 - k )^2"

"9-6h+h^2+9-6k+k^2""=1-2h+h^2+25-10k+k^2"

"18-6h-6k=26-2h-10k"

"4h=-8+4k"

"h=-2+k"

Substitute


"-2+k=11-2k"

"k=\dfrac{13}{3}"

"h=\dfrac{7}{3}"

"( 3 - \dfrac{7}{3} )^2 + (3 - \dfrac{13}{3} )^2 = r^2"

"r^2=\dfrac{20}{9}"

The center of the circle is "C( \dfrac{7}{3},\dfrac{13}{3})." The radius of the circle is "\dfrac{2\sqrt{5}}{3}."

The general equation of the circle is 


"(x-\dfrac{7}{3})^2 + ( y-\dfrac{13}{3} )^2 = \dfrac{20}{9}"